# Enhancing Language Multi-Agent Learning with Multi-Agent Credit Re-Assignment for Interactive Environment Generalization

### v1

@@ Abstract
LLM-based agents have made significant advancements in interactive environments, such as mobile operations and web browsing, and other domains beyond computer using. Current multi-agent systems universally excel in performance, compared to single agents, but struggle with generalization across environments due to predefined roles and inadequate strategies for generalizing language agents. The challenge of achieving both strong performance and good generalization has hindered the progress of multi-agent systems for interactive environments. To address these issues, we propose CollabUIAgents, a multi-agent reinforcement learning framework with a novel multi-agent credit re-assignment (CR) strategy, assigning process rewards with LLMs rather than environment-specific rewards and learning with synthesized preference data, in order to foster generalizable, collaborative behaviors among the role-free agents’ policies. Empirical results show that our framework improves both performance and cross-environment generalizability of multi-agent systems. Moreover, our 7B-parameter system achieves results on par with or exceed strong closed-source models, and the LLM that guides the CR. We also provide insights in using granular CR rewards effectively for environment generalization, and accommodating trained LLMs in multi-agent systems. Our data and code are available at https://github.com/THUNLP-MT/CollabUIAgents.
Enhancing Language Multi-Agent Learning with Multi-Agent
Credit Re-Assignment for Interactive Environment Generalization
Zhitao He3*Zijun Liu2*, Peng Li1, May Fung3, Ming Yan4 Ji Zhang4, Fei Huang4, Yang Liu1,2 1Institute for AI Industry Research (AIR), Tsinghua University 2Department of Computer Science and Technology, Tsinghua University 3Hong Kong University of Science and Technology, 4Alibaba Group zhitao.he@connect.ust.hk, zj-liu24@mails.tsinghua.edu.cn
@@ 1 Introduction
Autonomous agents have made substantial progress in interactive environments, such as mobile operations and web browsing, by leveraging large language models (LLMs). These agents hold immense potential not only to automate repetitive tasks but also to enhance decision-making and streamline complex workflows. As a result, they can free up human resources for higher-level problem-solving and innovation. The increasing interest in developing such agents is evident in the growing body of work on, for instance, mobile (Rawles et al., 2023, 2025; Zhang et al., 2024b; Deng et al., 2024a; Wang et al., 2024c), web browsing (Shi et al., 2017; Liu et al., 2018a; Yao et al., 2022; Zhou et al., 2024a; Deng et al., 2023, 2024b), and computer using environments (Xie et al., 2024; Sun et al., 2024), and LLM-based agents targeting on these tasks, including single-agent (Yan et al., 2023; Wang et al., 2024b; Hong et al., 2024b; Cheng et al., 2024a; Hu et al., 2024; Zhang et al., 2025) and multi-agent systems (Wang et al., 2024a; Zhou et al., 2023; Liu et al., 2024; Zhang et al., 2024c).
@@
In this work, we introduce a reinforcement learning framework for language multi-agent systems, named CollabUIAgents, designed to address challenges above in real-world interactive environments. The methodology is inspired from the success of classic multi-agent reinforcement learning (MARL) to stimulate collaboration (Ma and Luo, 2022). Beyond previous work (Tran et al., 2025), collaboration among langugae agents may also be beneficial for generalization across environments, as illustrated in Figure 1. Compared to existing methods using LLMs in credit assignment for non-language multi-agent settings (Qu et al., 2025; Lin et al., 2025), further considerations are made in our framework to enrich sparse rewarding signals, enhance generalization, and mitigate errors in credit assignment. We propose a novel multi-agent credit re-assignment (CR) strategy, assigning process rewards without using environment-specific outcome rewards, but with the world knowledge embedded in LLMs, and learning with synthesized preference data, aiming to foster generalizable, collaborative behaviors among the role-free agents’ policies.
@@ 2 Formulation and Notations
Interactive Environment  The observation and action space in interactive environments are enormous. Specifically, for the mobile operation environments, which offer an interface that allows agents to receive observations and perform actions on mobile devices, the observation space may include high-resolution screenshots and a UI tree from Android automater. The action space mirrors human interactions, featuring gestures (such as tapping, long-pressing, and swiping), typing, and navigation buttons (e.g., home and back). Complete actions are listed in Table 5. For web browsing environments, the observation space may include task description, simplified HTML, and current location. The HTML offers the model both structural and content details of the page, while the current location information allows it to understand its position on the webpage. Consistent with previous work (Lai et al., 2024), we use a unified web browsing action space in both of the aforementioned environments. The actions include hover, select, click, etc. Complete actions are listed in Table 6.
@@
Figure 2: The CollabUIAgents framework with credit re-assignment. The critic agent assess the environment state and the action matrix to get granular rewards. Agents are optimized with synthesized preference data from the adversarial agent to learn rewarded behaviors.
@@ 3.1 Multi-Agent System Architecture
The agents operate in a topological order, and starting from the source to the sink node, allowing each agent to aggregate all responses from its predecessors to form $C^{i,j}$ in Equation 2. We define the round of conversation as $m$. In each conversation round, all agents output an action and messages once along the DAG, and each agent receives outputs from itself at last round besides from predecessors, i.e., we keep a contextual memory of the size equal to 1. The proper size of contextual memories enhances the diversity of decision making and avoids introducing too long contexts during multi-agent communications. According to Equation 2, at the time step $t$ to interact with the environment, the system produce an action matrix ${\bm{A}}_{t}=(a_{t}^{i,j}),i=1,...,n;j=1,...,m$. Then, majority voting is used to decide the final action
@@ 3.2 Language Multi-Agent RL
Credit Re-Assignment  Traditional MARL systems use value decomposition for credit assignment (Ma and Luo, 2022), and recent work also uses multiple value estimators to alleviate reward sparcity in interactive environments (Bai et al., 2024). However, training critic models for each agent based on outcome rewards is computationally expensive and may hinder generalization. Instead, by using an LLM-based critic agent, represented by $\pi_{\mathrm{critic}}$, we provide process rewards in a finer granularity for each agent at each conversation round. The action matrix ${\bm{A}}_{t}$ is assessed by the critic agent based on observed environment state $o_{t}$, interaction history $H_{t-1}$, and the task query $q$, generating a reward matrix ${\bm{R}}_{t}=(r_{t}^{i,j})$, $i=1,...,n,j=1,...,m$, where $r_{t}^{i,j}\in\{0,1\}$ denotes the intermediate reward for $a_{t}^{i,j}$:
@@
$\mathcal{L}_{\text{MARL}}(\pi_{i})=-\mathbb{E}_{\mathcal{E}_{\mathcal{G}}^{\prime}\sim K_{|\mathcal{G}|}}\mathbb{E}_{(s_{t},a_{t}^{i,j},\hat{H}_{t}^{i})\sim\mathcal{P}(\mathcal{G},\mathcal{E}_{\mathcal{G}}^{\prime})}$ (8)
$\sum_{t=0}^{T_{\mathrm{max}}}\sum_{j=1}^{m}\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta_{i}}(a_{t}^{i,j}|o_{t},\hat{H}_{t}^{i})}{\log\pi_{\mathrm{ref}_{i}}(a_{t}^{i,j}|o_{t},\hat{H}_{t}^{i})}\right.\right.\right.$
$\left.\left.\left.-\frac{\log\pi_{\theta_{i}}(a_{t}^{i,j,-}|o_{t},\hat{H}_{t}^{i})}{\log\pi_{\mathrm{ref}_{i}}(a_{t}^{i,j,-}|o_{t},\hat{H}_{t}^{i})}\right)\right)\right]\cdot\bm{1}_{r_{t}^{i,j}=1},$
@@ 3.3 Cross-Environment Adaptation
Method Base model #Params #Agents Input SRAndroidWorld SRMMiniWoB++ $\bm{\Delta}$Generalization
@@
Table 1: Experimental results on mobile operation environments. Success rates (SR) in AndoridWorld and MobileMiniWoB++ (MMiniWoB++) are listed. $\bm{\Delta}$Generalization indicates the performance gap between the base model and agent learning methods based on the model in MobileMiniWoB++.
@@ 4.1 Experimental Settings
Environments  We conduct experiments in both mobile operation and web browsing environments. For the mobile environments, we use AndroidWorld (Rawles et al., 2025) for training and MobileMiniWoB++ (Rawles et al., 2025) for testing: (1) AndroidWorld has 116 programmatic tasks across 20 real-world apps, such as Chrome, Markor, and Pro Expense. (2) MobileMiniWoB++ is derived from MiniWoB++ (Shi et al., 2017), which is a web-based benchmark. MobileMiniWoB++ shares the same observation space as AndroidWorld and supports 92 tasks from MiniWoB++. We use the success rate (SR) as an evaluation metric. For the web environments, we leverage Mind2Web (Deng et al., 2023) for training and AutoWebBench (Lai et al., 2024) for testing: (1) Mind2Web features over 2,000 open-ended tasks sourced from 137 websites in 31 different domains. (2) AutoWebBench is a bilingual benchmark featuring approximately 10,000 traces, from mainstream Chinese and English websites, providing a diverse dataset for web browsing. We use the step-success rate (SSR) as the evaluation metric. For agent learning methods on open-source LLMs, we use the performance gap between the base model and the trained model in unseen environments, $\bm{\Delta}$Generalization, to indicate generalizability.
@@
Method Base model #Params #Agents Input SSRMind2Web SSRAutoWebBench $\bm{\Delta}$Generalization
@@
Base Model LLaMA2 7B 1 Text 2.9 -
@@
Table 2: Experimental results on web browsing environments. Average step success rates (SSR) in Mind2Web and AutoWebBench are reported. “SFT” denotes the base model is supervised fine-tuned in Mind2Web. $\bm{\Delta}$Generalization indicates the gap between the base model and agent learning methods based on the model in AutoWebBench.
@@ 5 Related Work
Agents on Interactive Environments  Before the advent of LLMs, agents relied on traditional RL to perform interactions such as clicking and typing (Liu et al., 2018b; Humphreys et al., 2022). However, recent advancements have shifted towards leveraging foundation models with in-context learning or fine-tuning across various interfaces, including mobile (Wang et al., 2023; Hong et al., 2024b), web (Lai et al., 2024; Deng et al., 2024b), and computer using environments (Xu et al., 2024; Wu et al., 2024b). Recently, there are emerging methods designing process rewards (He et al., 2024; Pan et al., 2024) or language RL (Bai et al., 2024) for better performing single agents.
Interactive Environments for Agents  To effectively evaluate language agents, it is essential to create environments that replicate real-world conditions and deliver accurate rewards (Rawles et al., 2023; Deng et al., 2023). MiniWoB++ (Shi et al., 2017) is a lightweight framework that features small, synthetic HTML pages with parameterized tasks. WebArena (Zhou et al., 2024b) and its visual counterpart, VisualWebArena (Koh et al., 2024), simulate websites spanning up to distinct domains, while WorkArena (Drouin et al., 2024) focuses on enterprise software. For more specialized environments, WebShop (Yao et al., 2022) simulates an e-commerce platform for online shopping. For computer using environment, OSWorld (Xie et al., 2024) provides both a user interface and programmatically generated rewards across different apps.
Prompt-Based Multi-agent Learning  Collaboration among multiple LLM agents has shown effective for various tasks (Hong et al., 2024a; Wu et al., 2024a; Qian et al., 2024; Wang et al., 2025). However, employing a static architecture without team optimization may restrict the performance and generalization. Chen et al. (2024) selects a fixed number of agents from a set of manual prompt candidates via an additional LLM during each round of discussion. Zhuge et al. (2024b) unify language agent systems by describing them as optimizable computational graphs and develop optimization methods for nodes and edges, enabling automatic improvements of agent prompts and inter-agent orchestration. Liu et al. (2024) employ a layered network to formulate the process of LLM-agent collaboration for arbitrary tasks and introduce an unsupervised algorithm to optimize the team of agents by the individual contributions of agent.
@@ 6 Conclusion
Limitations
Current methodology of CollabUIAgents framework achieves significant improvement in both environment-specific performance and cross-environment generalization. Howevrer, it still has several limitations. First, the proposed multi-agent system is based on a fixed number of agents, which may not be optimal for all tasks. Second, the rolling-out phase in the MARL process relies on majority voting, which may limit the diversity of the rolled-out actions and could be further optimized. Third, Third, due to computational resource limitations, we have not explored large-scale MARL training, such as with more agents, larger query sets, or simply more training steps, compared to large-scale RL practices (DeepSeek-AI, 2025). The potential of reinforcement learning on language agents deserves further efforts.
@@ B.1 Methodology
(2) Simple Instruction Knowledge: The agents are tasked with performing basic interactions, such as clicking or typing, in response to simple instructions. Specifically, given the complete action space, we prompt the UI agent to generate possible instructions related to a random UI element, and their corresponding responses. For example, in Figure 3, the UIAgent was prompted to generate an instruction for element 9 (“selecting the M4a format") and then generates the corresponding response to interact with it. By learning this type of knowledge, the agent lays the foundation for completing a complex sequential decision-making process.
@@
$\mathcal{L}_{\text{DPO}}=$ $-\mathbb{E}_{(s,a^{+},a^{-})\sim\mathcal{P}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(a^{-}|s)}{\pi_{\mathrm{ref}}(a^{-}|s)}\right.\right.$ (10)
$\left.\left.-\beta\log\frac{\pi_{\theta}(a^{+}|s)}{\pi_{\mathrm{ref}}(a^{+}|s)}\right)\right],$

### v2

# Advancing Language Multi-Agent Learning with Credit Re-Assignment for Interactive Environment Generalization

@@ Abstract
LLM-based agents have made significant advancements in interactive environments, such as mobile operations and web browsing, and other domains beyond computer using. Current multi-agent systems universally excel in performance, compared to single agents, but struggle with generalization across environments due to predefined roles and inadequate strategies for generalizing language agents. The challenge of achieving both strong performance and good generalization has hindered the progress of multi-agent systems for interactive environments. To address these issues, we propose CollabUIAgents, a multi-agent reinforcement learning framework with a novel multi-agent credit re-assignment (CR) strategy, assigning process rewards with LLMs rather than environment-specific rewards and learning with synthesized preference data, in order to foster generalizable, collaborative behaviors among the role-free agents’ policies. Empirical results show that our framework improves both performance and cross-environment generalizability of multi-agent systems. Moreover, our 7B-parameter system achieves results on par with or exceed strong closed-source models, and the LLM that guides the CR. We also provide insights in using granular CR rewards effectively for environment generalization, and accommodating trained LLMs in multi-agent systems. Our work is available at https://github.com/THUNLP-MT/CollabUIAgents.
@@ 1 Introduction
Autonomous agents have made substantial progress in interactive environments, such as mobile operations and web browsing, by leveraging large language models (LLMs). These agents hold immense potential not only to automate repetitive tasks but also to enhance decision-making and streamline complex workflows. As a result, they can free up human resources for higher-level problem-solving and innovation. The increasing interest in developing such agents is evident in the growing body of work on, for instance, mobile (Rawles et al., 2023; 2025; Zhang et al., 2024b; Deng et al., 2024a; Wang et al., 2024c), web browsing (Shi et al., 2017; Liu et al., 2018a; Yao et al., 2022; Zhou et al., 2024b; Deng et al., 2023; 2024b), and computer using environments (Xie et al., 2024; Sun et al., 2024), and LLM-based agents targeting on these tasks, including single-agent (Yan et al., 2023; Wang et al., 2024b; Hong et al., 2024b; Cheng et al., 2024a; Hu et al., 2024; Zhang et al., 2025) and multi-agent systems (Wang et al., 2024a; Zhou et al., 2023; Zhang et al., 2024c).
@@
In this work, we introduce a reinforcement learning framework for language multi-agent systems, named CollabUIAgents, designed to address challenges above in real-world interactive environments. The methodology is inspired from the success of classic multi-agent reinforcement learning (MARL) to stimulate collaboration (Ma & Luo, 2022). Beyond previous work (Tran et al., 2025), collaboration among langugae agents may also be beneficial for generalization across environments, as illustrated in Figure 1. Compared to existing methods using LLMs in credit assignment for non-language multi-agent settings (Qu et al., 2025; Lin et al., 2025), further considerations are made in our framework to enrich sparse rewarding signals, enhance generalization, and mitigate errors in credit assignment. We propose a novel multi-agent credit re-assignment (CR) strategy, assigning process rewards without using environment-specific outcome rewards, but with the world knowledge embedded in LLMs, and learning with synthesized preference data, aiming to foster generalizable, collaborative behaviors among the role-free agents’ policies.
@@ 2 Formulation and Notations
Interactive Environment  The observation and action space in interactive environments are enormous. Specifically, for the mobile operation environments, which offer an interface that allows agents to receive observations and perform actions on mobile devices, the observation space may include high-resolution screenshots and a UI tree from Android automater. The action space mirrors human interactions, featuring gestures (such as tapping, long-pressing, and swiping), typing, and navigation buttons (e.g., home and back). Complete actions are listed in Table 5. For web browsing environments, the observation space may include task description, simplified HTML, and current location. The HTML offers the model both structural and content details of the page, while the current location information allows it to understand its position on the webpage. Consistent with previous work (Lai et al., 2024), we use a unified web browsing action space in both of the aforementioned environments. The actions include hover, select, click, etc. Complete actions are in Table 6.
@@
Figure 2: The CollabUIAgents framework with credit re-assignment. The critic agent assess the environment state and the action matrix to get granular rewards. Agents are optimized with synthesized preference data to learn rewarded behaviors.
@@ 3.1 Multi-Agent System Architecture
The agents operate in a topological order, and starting from the source to the sink node, allowing each agent to aggregate all responses from its predecessors to form $C^{i,j}$ in Equation 2. We define the round of conversation as $m$. In each conversation round, all agents output an action and messages once along the DAG, and each agent receives outputs from itself at last round besides from predecessors, i.e., we keep a contextual memory of the size equal to 1. The proper size of contextual memories enhances the diversity of decision making and avoids introducing too long contexts during multi-agent communications. According to Equation 2, at the time step $t$, the system produce an action matrix $\bm{A}_{t}=(a_{t}^{i,j}),i=1,...,n;j=1,...,m$. Then, majority voting is used to decide the final action:
@@ 3.2 Language Multi-Agent RL
Credit Re-Assignment  Traditional MARL systems use value decomposition for credit assignment (Ma & Luo, 2022), and recent work also uses multiple value estimators to alleviate reward sparcity in interactive environments (Bai et al., 2024). However, training critic models for each agent based on outcome rewards is computationally expensive and may hinder generalization. Instead, by using an LLM-based critic agent, represented by $\pi_{\mathrm{critic}}$, we provide process rewards in a finer granularity for each agent at each conversation round. The action matrix $\bm{A}_{t}$ is assessed by the critic agent based on observed environment state $o_{t}$, interaction history $H_{t-1}$, and the task query $q$, generating a reward matrix $\bm{R}_{t}=(r_{t}^{i,j})$, $i=1,...,n,j=1,...,m$, where $r_{t}^{i,j}\in\{0,1\}$ denotes the intermediate reward for $a_{t}^{i,j}$:
@@
$\mathcal{L}_{\text{MARL}}(\pi_{i})=-\mathbb{E}_{\mathcal{E}_{\mathcal{G}}^{\prime}\sim K_{|\mathcal{G}|}}\mathbb{E}_{(s_{t},a_{t}^{i,j},\hat{H}_{t}^{i})\sim\mathcal{P}(\mathcal{G},\mathcal{E}_{\mathcal{G}}^{\prime})}\\ \sum_{t=0}^{T_{\mathrm{max}}}\sum_{j=1}^{m}\left[\log\sigma\left(\beta\left(\frac{\log\pi_{\theta_{i}}(a_{t}^{i,j}|o_{t},\hat{H}_{t}^{i})}{\log\pi_{\mathrm{ref}_{i}}(a_{t}^{i,j}|o_{t},\hat{H}_{t}^{i})}\right.\right.\right.\left.\left.\left.-\frac{\log\pi_{\theta_{i}}(a_{t}^{i,j,-}|o_{t},\hat{H}_{t}^{i})}{\log\pi_{\mathrm{ref}_{i}}(a_{t}^{i,j,-}|o_{t},\hat{H}_{t}^{i})}\right)\right)\right]\cdot\bm{1}_{r_{t}^{i,j}=1},$ (8)
@@ 3.3 Cross-Environment Adaptation
Method Base model #Params #Agents Input SRAndroidWorld SRMMiniWoB++ $\boldsymbol{\Delta}$Generalization
@@
Table 1: Experimental results on mobile operation environments. Success rates (SR) in AndoridWorld and MobileMiniWoB++ (MMiniWoB++) are listed. $\boldsymbol{\Delta}$Generalization indicates the performance gap between the base model and agent learning methods based on the model in MobileMiniWoB++.
@@ 4.1 Experimental Settings
Environments  We conduct experiments in both mobile operation and web browsing environments. For the mobile environments, we use AndroidWorld (Rawles et al., 2025) for training and MobileMiniWoB++ (Rawles et al., 2025) for testing: (1) AndroidWorld has 116 programmatic tasks across 20 real-world apps, such as Chrome, Markor, and Pro Expense. (2) MobileMiniWoB++ is derived from MiniWoB++ (Shi et al., 2017), which is a web-based benchmark. MobileMiniWoB++ shares the same observation space as AndroidWorld and supports 92 tasks from MiniWoB++. We use the success rate (SR) as an evaluation metric. For the web environments, we leverage Mind2Web (Deng et al., 2023) for training and AutoWebBench (Lai et al., 2024) for testing: (1) Mind2Web features over 2,000 open-ended tasks sourced from 137 websites in 31 different domains. (2) AutoWebBench is a bilingual benchmark featuring approximately 10,000 traces, from mainstream Chinese and English websites, providing a diverse dataset for web browsing. We use the step-success rate (SSR) as the evaluation metric. For agent learning methods on open-source LLMs, we use the performance gap between the base model and the trained model in unseen environments, $\boldsymbol{\Delta}$Generalization, to indicate generalizability.
@@
Method Base model #Params #Agents Input SSRMind2Web SSRAutoWebBench $\boldsymbol{\Delta}$Generalization
@@
Base Model LLaMA2 7B 1 Text - 2.9 -
@@
Table 2: Experimental results on web browsing environments. Average step success rates (SSR) in Mind2Web and AutoWebBench are reported. “SFT” denotes the base model is supervised fine-tuned in Mind2Web. $\boldsymbol{\Delta}$Generalization indicates the gap between the base model and agent learning methods based on the model in AutoWebBench.
@@ 5 Related Work
Agents on Interactive Environments  Before the advent of LLMs, agents relied on traditional RL to perform interactions such as clicking and typing (Liu et al., 2018b; Humphreys et al., 2022). However, recent advancements have shifted towards leveraging foundation models with in-context learning or fine-tuning across various interfaces, including mobile (Wang et al., 2023; Hong et al., 2024b), web (Lai et al., 2024; Deng et al., 2024b), and computer using environments (Xu et al., 2024; Wu et al., 2024b). Recently, there are emerging methods designing process rewards (He et al., 2024a; Pan et al., 2024; He et al., 2025; Huang et al., 2025) or language RL (Bai et al., 2024) for better performing single agents.
Interactive Environments for Agents  To effectively evaluate language agents, it is essential to create environments that replicate real-world conditions and deliver accurate rewards (Rawles et al., 2023; Deng et al., 2023). MiniWoB++ (Shi et al., 2017) is a lightweight framework that features small, synthetic HTML pages with parameterized tasks. WebArena (Zhou et al., 2024a) and its visual counterpart, VisualWebArena (Koh et al., 2024), simulate websites spanning up to distinct domains, while WorkArena (Drouin et al., 2024) focuses on enterprise software. For more specialized environments, WebShop (Yao et al., 2022) simulates an e-commerce platform for online shopping.
Prompt-Based Multi-agent Learning  Collaboration among multiple LLM agents has shown effective for various tasks (He et al., 2023; Hong et al., 2024a; Wu et al., 2024a; He et al., 2024c; Jin et al., 2024; Qian et al., 2024; He et al., 2024b; Wang et al., 2025). However, employing a static architecture without team optimization may restrict the performance and generalization. Chen et al. (2024) selects a fixed number of agents from a set of manual prompt candidates via an additional LLM during each round of discussion. Zhuge et al. (2024b) unify language agent systems by describing them as optimizable computational graphs and develop optimization methods for nodes and edges, enabling automatic improvements of agent prompts and inter-agent orchestration.
@@ B.1 Methodology
(2) Simple Instruction Knowledge: The agents are tasked with performing basic interactions, such as clicking or typing, in response to simple instructions. Specifically, given the complete action space, we prompt the UI agent to generate possible instructions related to a random UI element, and their corresponding responses. For example, in Figure 3, the UIAgent was prompted to generate an instruction for element 9 (“selecting the M4a format”) and then generates the corresponding response to interact with it. By learning this type of knowledge, the agent lays the foundation for completing a complex sequential decision-making process.
@@
$\mathcal{L}_{\text{DPO}}=$ $-\mathbb{E}_{(s,a^{+},a^{-})\sim\mathcal{P}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(a^{-}|s)}{\pi_{\mathrm{ref}}(a^{-}|s)}\right.\right.\left.\left.-\beta\log\frac{\pi_{\theta}(a^{+}|s)}{\pi_{\mathrm{ref}}(a^{+}|s)}\right)\right],$ (10)